# Beyond Autoregression: Fast LLMs via Self-Distillation Through Time

### v1

@@ 4.2 Additional ablations
Figure 4: MAUVE performance. MAUVE performance of the student after distillation. The sub-title dt=1/k refers to the decoding step size that the student should use to match the teacher. For example, the plot with title dt=1/256 showcases the MAUVE score after two rounds of SDTT. The teacher performance is computed from the original, non-distilled teacher using 128 decoding steps. The dashed line denotes the best MAUVE score achieved by the students being compared.
@@
In section 4.1, each round of SDTT consists of $10k$ steps. Since the magnitude of the distillation loss does not reliably indicate convergence, we experiment with shorter rounds. We find that reducing the number of steps to $5k$ or $2.5k$ negatively impacted conditional generation performance, as shown in fig. 6. However, shorter rounds slightly improved the final generative perplexity (fig. 7) and resulted in marginally better LAMBADA perplexity (fig. 8). Since SDTT does not directly optimize the ELBO, an increase in perplexity is expected. Interestingly, the LAMBADA accuracy remains unchanged with shorter rounds.
@@
In section 4.1, the targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ are generated using 2 sampling steps from the teacher. We explore distilling a larger number of sampling steps at once (4 or 8), since using more rounds of SDTT may induce more error accumulation in approximating the original teacher. Figure 11 shows that distilling more than two steps at a time is difficult and results in weaker results on LAMBADA. This result suggests that the higher stochasticity of the targets generated with four or eight steps makes the task too difficult for the student.
@@
Lou et al. (2023) observe that using an analytical sampler (Campbell et al., 2022) results in higher quality samples compared to ancestral sampling. However, when generating targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ with analytical sampling, we observed minimal difference with ancestral sampling, as shown in fig. 9 and 10.
@@
Using an Exponential Moving Average (EMA) of the weights is known to improve the quality of samples from diffusion models (Nichol & Dhariwal, 2021). However, when applying SDTT for multiple rounds, it is unclear whether the EMA or current weights should be used as the teacher for successive rounds. Additionally, it could be favorable to reset the optimizer state between rounds as we grow the decoding step size. We experiment with two approaches: either resetting the optimizer state only, or resetting both the EMA and optimizer state. Figure 12 shows the generative perplexity when resetting the optimizer state and using the EMA as the teacher instead of the current weights, while fig. 13 presents the corresponding results for MAUVE. When using the EMA as teacher, since we accumulate updates in the EMA over 10k steps only, we use a slightly lower decay rate of 0.999. The results indicate that using the EMA of the weights as the teacher may slightly improve performance.
@@ 4.3 Scaling SDTT to 860M parameters
Figure 5: Perplexity for different losses and decoding step size. Generative perplexity over 7 rounds of SDTT with MSE, TVD and KLD. While the KLD leads to a higher perplexity than the MSE; we focus on the KLD because it is the only divergence that retains the performance on the LAMBADA dataset. The curve labeled as dt=1/k denotes the student trained to match the teacher distribution with $k$ decoding steps. For example, the performance of the student after 4 rounds of SDTT is labeled as dt=1/128 as we double the decoding step size after each round of SDTT.
We apply SDTT to larger discrete diffusion models with up to 860M parameters. In this experiment, we train the models from scratch for 400k steps with a batch size of 512, a context length of 1024 and the Adam optimizer. We reuse the training configuration of Sahoo et al. (2024) and scale the models to larger sizes. We train 3 model sizes, small (169M), medium (424M) and large (863M). Details of the model architecture for each scale are shown in table 1. As in the other experiments, the models are diffusion transformers (Peebles & Xie, 2023) and we use an EMA with a decay of 0.9999. Although the results in section 4.2 suggest that short distillation rounds might be sufficient, it is unclear whether this result also holds on larger scales. Therefore, we use $10k$ steps per round of SDTT. For simplicity, we generate targets using 2 teacher ancestral decoding steps and do not reset the optimizer state or EMA between rounds.
@@ 4.4 Latency with SDTT
While SDTT allows sampling from discrete diffusion models with 32-64 times less decoding steps, a quantity of interest to practitioners is the actual latency of text generation. Indeed, while the reduction in the number of sampling steps is large, since discrete diffusion uses a non-causal architecture, we cannot use KV caching (Pope et al., 2022). KV caching improves the inference performance drastically for AR models, hence we compare the latency of SDTT with GPT-2 with KV caching. We successfully reproduce the results of Deschenaux & Gulcehre (2024), which showed a 4x improvement when sampling with 32 steps, only this time, we retain the text quality because of SDTT. We additionally measure the latency when decoding with 16 steps and observe an 8x speedup against the AR baseline. We compute the latency using untrained models with around 1.3B parameters, using the same hyperparameters as Deschenaux & Gulcehre (2024). We use a batch size of 8 and time the sampling 10 times after one warm-up step on a single A100 GPU with 80 GiB of RAM.
[…]
@@ 5 Related Work
Diffusion models Sohl-Dickstein et al. (2015b); Ho et al. (2020); Song & Ermon (2020) are the basis of many state-of-the-art text-to-image models (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022). After their introduction by Sohl-Dickstein et al. (2015b), Ho et al. (2020) showed that diffusion models can achieve FID (Heusel et al., 2017) scores comparable to GANs Goodfellow et al. (2014); Arjovsky et al. (2017).
@@
Prior to Sahoo et al. (2024); Shi et al. (2024); Ou et al. (2024), Lou et al. (2023) introduced a novel discrete diffusion language model called SEDD. When decoding with a large number of steps, SEDD can match or surpass GPT-2 in unconditional text generation. The model of Lou et al. (2023) learn a discrete generalization of the score of continuous diffusion models (Song & Ermon, 2020; Song et al., 2021). Campbell et al. (2022); Zhao et al. (2024) developed the continuous-time discrete diffusion framework. Hoogeboom et al. (2021) extended Bernoulli diffusion (Sohl-Dickstein et al., 2015b) to categorical distributions, and Austin et al. (2023) generalized the work of Hoogeboom et al. (2021) to more general corruption processes, including absorbing diffusion, a key component of later work. In parallel, several studies have explored continuous diffusion for language modeling (Li et al., 2022; Dieleman et al., 2022; Han et al., 2023; Chen et al., 2023; Gulrajani & Hashimoto, 2024). Despite recent breakthroughs, diffusion language models still have some drawbacks (Deschenaux & Gulcehre, 2024).
@@ 6 Discussion
In this work, we introduce Self-Distillation Through Time (SDTT), a distillation method for discrete diffusion models. Recent work (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2024) suggests that discrete diffusion models can match or outperform autoregressive models in text quality. However, those models require more inference resources than AR models to achieve good performance, because of the non-causal architecture of the neural network that prevents the use of KV caching. We show that SDTT can reduce the number of decoding steps while retaining performance. Our final student is up to 8x faster than AR models that use KV caching and we demonstrate that SDTT is applicable to larger models as well. In future work, we plan to evaluate SDTT on tasks that involve generating a large number of completions from a base language model.
@@ 7 Reproducibility Statement
We provide details on model architectures, hyperparameters, and provide pseudocode for our algorithm. We built on top of the open source model of Sahoo et al. (2024), which makes it relatively easy for researchers to reproduce our results. Additionally, we release code and checkpoints on GitHub and HuggingFace.
@@ Appendix A Additional ablation results
(a) Resetting the optimizer state between rounds.
@@ Appendix B Additional details on the divergence measures
Model size small medium large
# params 169M 424M 863M
Num Layers 12 24 24
Embedding dim. 768 1024 1536
Num. heads 12 16 16
Table 1: Hyperparameters of the diffusion models at different scales. All models have a context length of 1024 tokens and use RoPE positional encoding (Su et al., 2023).
@@ Appendix D Text examples
We include non-cherry picked text generated from the small distilled model with KLD loss from the last round of distillation via unconditional sampling with varying number of steps. We show the first 512 tokens to so that the text fits on one page. Remember that those models small and not fine-tuned for text quality. They also typically start generating in the middle of sentences, since they are trained on a concatenated corpus of documents.

### v2

@@ 4.2 Additional ablations
Figure 5: MAUVE performance of the student after each round of SDTT. The teacher performance is computed using samples generated with 128 decoding steps.
@@
In section 4.1, each round of SDTT consists of $10k$ training iterations. Since the magnitude of the distillation loss does not reliably indicate convergence, we experiment with shorter rounds. We find that reducing the number of training iterations to $5k$ or $2.5k$ negatively impacted conditional generation performance, as shown in fig. 7. However, shorter rounds slightly improved the final generative perplexity (fig. 8) and resulted in marginally better LAMBADA perplexity (fig. 10). Since SDTT does not directly optimize the ELBO, an increase in perplexity is expected. Interestingly, the LAMBADA accuracy remains unchanged with shorter rounds.
@@
In section 4.1, the targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ are generated using 2 sampling steps from the teacher. We explore distilling a larger number of sampling steps at once (4 or 8), since using more rounds of SDTT may induce more error accumulation in approximating the original teacher. Figure 13 shows that distilling more than two steps at a time is difficult and results in weaker results on LAMBADA. This suggests that the higher stochasticity of the targets generated with four or eight steps makes the task too difficult for the student.
@@
Lou et al. (2023) observe that using an analytical sampler (Campbell et al., 2022) results in higher quality samples compared to ancestral sampling. However, when generating targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ with analytical sampling, we observed minimal difference with ancestral sampling, as shown in fig. 11 and 12.
@@
Using an Exponential Moving Average (EMA) of the weights is known to improve the quality of samples from diffusion models (Nichol & Dhariwal, 2021). However, when applying SDTT for multiple rounds, it is unclear whether the EMA or current weights should be used as the teacher for successive rounds. Additionally, it could be favorable to reset the optimizer state between rounds as we grow the decoding step size. We experiment with two approaches: either resetting the optimizer state only, or resetting both the EMA and optimizer state. Figure 14 shows the generative perplexity when resetting the optimizer state and using the EMA as the teacher instead of the current weights, while fig. 15 presents the corresponding results for MAUVE. When using the EMA as teacher, since we accumulate updates in the EMA over 10k training iterations only, we use a slightly lower decay rate of 0.999. We find that using the EMA of the weights as the teacher may slightly improve performance.
@@ 4.3 Scaling SDTT to 860M parameters
Figure 6: Perplexity for different losses and decoding step size. Generative perplexity over 7 rounds of SDTT with MSE, TVD and KLD. While the KLD leads to a higher perplexity than the MSE; we focus on the KLD because it is the only divergence that retains the performance on the LAMBADA dataset.
We apply SDTT to larger discrete diffusion models with up to 860M parameters. In this experiment, we train the models from scratch for 400k steps with a batch size of 512, a context length of 1024 and the Adam optimizer. We reuse the training configuration of Sahoo et al. (2024) and scale the models to larger sizes. We train 3 model sizes, small (169M), medium (424M) and large (863M). Details of the model architecture for each scale are shown in table 2. As for the other experiments, the models are diffusion transformers (Peebles & Xie, 2023) and we use an EMA with a decay of 0.9999. Although the results in section 4.2 suggest that short distillation rounds might be sufficient, it is unclear whether this result also holds on larger scales. Therefore, we use $10k$ steps per round of SDTT. For simplicity, we generate targets using 2 teacher ancestral decoding steps and do not reset the optimizer state or EMA between rounds.
@@ 4.4 Latency with SDTT
While SDTT allows sampling from discrete diffusion models with 32-64 times less decoding steps, a quantity of interest to practitioners is the actual latency of text generation. Indeed, while the reduction in the number of sampling steps is large, since discrete diffusion uses a non-causal architecture, we cannot use KV caching (Pope et al., 2022). KV caching improves the inference performance drastically for AR models, hence we compare the latency of SDTT with GPT-2 with KV caching. We successfully reproduce the results of Deschenaux & Gulcehre (2024), which showed a 4x improvement when sampling with 32 steps, and measure an 8x improvement with 16 decoding steps. We compute the latency using untrained models with around 1.3B parameters, using the same hyperparameters as Deschenaux & Gulcehre (2024). We use a batch size of 8 and time the sampling 10 times after one warm-up step on a single A100 GPU with 80 GiB of RAM. All models use FlashAttention (Dao et al., 2022). See Table 1 for additional experiments on the latency.
@@ 5 Related Work
Diffusion models (Sohl-Dickstein et al., 2015b; Ho et al., 2020; Song & Ermon, 2020) are the basis of many state-of-the-art text-to-image models (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022). After their introduction by Sohl-Dickstein et al. (2015b), Ho et al. (2020) showed that diffusion models can achieve FID scores (Heusel et al., 2017) comparable to GANs (Goodfellow et al., 2014; Arjovsky et al., 2017).
@@
Prior to Sahoo et al. (2024); Shi et al. (2024); Ou et al. (2024), Lou et al. (2023) introduced a novel discrete diffusion language model called SEDD. When decoding with a large number of steps, SEDD can match or surpass GPT-2 in unconditional text generation. The model of Lou et al. (2023) learn a discrete generalization of the score of continuous diffusion models (Song & Ermon, 2020; Song et al., 2021). Campbell et al. (2022); Zhao et al. (2024) developed the continuous-time discrete diffusion framework. Hoogeboom et al. (2021) extended Bernoulli diffusion (Sohl-Dickstein et al., 2015b) to categorical distributions, and Austin et al. (2023) generalized the work of Hoogeboom et al. (2021) to more general corruption processes, including absorbing diffusion. Zheng et al. (2024b) develop a family of re-parameterized discrete diffusion models to enhance the training and decoding efficiency. In parallel, several studies have explored continuous diffusion for language modeling (Li et al., 2022; Dieleman et al., 2022; Han et al., 2023; Chen et al., 2023; Gulrajani & Hashimoto, 2024). Despite recent breakthroughs, diffusion language models still have some drawbacks (Deschenaux & Gulcehre, 2024). Ye et al. (2024) adapt Chain-of-Thought reasoning (Wei et al., 2023) to diffusion models.
@@ 6 Discussion
In this work, we introduce Self-Distillation Through Time (SDTT), a distillation method for discrete diffusion models. Recent works (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2024) suggest that discrete diffusion models can match or outperform autoregressive models in text quality. However, those models require more inference resources than AR models to achieve good performance, because of the non-causal architecture of the neural network that prevents the use of KV caching. We show that SDTT can reduce the number of decoding steps while retaining performance. Our final student is up to 8x faster than AR models that use KV caching and we demonstrate that SDTT is applicable to larger models as well. In future work, we plan to evaluate SDTT on tasks that involve generating a large number of completions from a base language model.
@@ 7 Reproducibility Statement
We provide details on model architectures, hyperparameters, and provide pseudocode for our algorithm. We built on top of the open source model of Sahoo et al. (2024), which makes it relatively easy for researchers to reproduce our results. Additionally, upon de-anonymization, we will release our code and artifacts.
@@ 8 Ethics Statement
9 Acknowledgements
We thank the ICLR’25 reviewers, area chairs, and organizers for their valuable feedback and support. We acknowledge the SCITAS team at EPFL for providing access to their beta cluster, and Karin Gétaz for her administrative assistance. This work was supported by the Swiss AI Initiative through a grant from the Swiss National Supercomputing Centre (CSCS), project ID a10 on Alps. Special thanks to Skander Moalla for providing a reproducible compute infrastructure code template.
@@ Appendix A Additional ablation results
Table 1: Downstream evaluation results. We report the accuracy of GPT-2, the teacher and students after 7 rounds of SDTT. Distillation seems to minimally affect the downstream performance.
Task GPT-2 Teacher KLD student MSE student TVD student
ARC-Easy 43.81 40.91 40.57 40.45 40.32
ARC-Challenge 19.03 21.08 20.73 19.28 20.05
HellaSwag 28.92 30.50 29.65 29.10 29.18
MathQA 21.21 21.78 21.47 22.28 21.84
PIQA 62.89 59.74 59.85 58.11 58.16
WinoGrande 51.62 50.91 50.75 49.57 50.36
@@
Generative perplexity and precision of the floating-point operations.
Zheng et al. (2024a) observed that low-precision sampling can be problematic in masked diffusion models, leading to reduced diversity and potentially misleading generative perplexity scores. As such, in addition to bfloat16, we try distilling (i.e. computing the backward KL) and sampling using 64 bits precision. Overall, it does lead to a higher generative perplexity, however the conclusions remain similar, as the final student achieves lower generative perplexity than GPT-2 with nucleus sampling (p=0.95) in 64 sampling steps, as shown in fig. 9.
@@
Conditional perplexity with TVD
In fig. 18(c), we show the conditional perplexity (prompt excluded) on the small scale, for models trained for 1M steps. Empirically, the TVD performs worse than the KLD and MSE.
Measuring the diversity
We evaluate the generation diversity using the self-BLEU score (Zhu et al., 2018). The self-BLEU score averages the BLEU score between one completion and the others. Therefore, when the sampling algorithm is deterministic, the self-BLEU score is 1, and a lower self-BLEU score denotes a more diverse set of samples. Formally, let $X=\{x_{1},...,x_{n}\}$ be conditionally-generated sequences, starting with the same prompt. The self-BLEU score can be computed as
[…]
We compute the self-BLEU score using 1000 prompts, as for MAUVE, and generate 5 continuations per prompt. Figure 4(a), fig. 19(a) and fig. 19(b) show the self-bleu score after distillation with the KLD, MSE and TVD objectives. Each objective only minimally decrease the diversity after distillation. Compared to on-policy distillation of autoregressive models (Agarwal et al., 2024), the decrease is marginal, as Agarwal et al. (2024) observe an increase of self-BLEU of the order of 10-20, demonstrating a more significant decrease in diversity.
Decoding latency
In addition to the results on the 1.3B scale, we report the latency for models with 169M, 424M, 863M, 3B and 8B parameters. We compute the latency with a batch size of 8 and 4. Figure 20 shows the latency with a batch size of 8 and fig. 21 using a batch size of 4. Figure 22 shows the trade-off between latency and perplexity. We measure the latency at the small model size and compare GPT-2 with the final students after 7 rounds of distillation.
Additional downstream evaluation results
We show the performance of GPT-2, the teacher and distilled students on additional downstream benchmarks from Gao et al. (2021) in table 1.
@@
Figure 9: Generative perplexity when distilling and sampling with 64 bits precision. Namely, we sample from the teacher and students in float64, and compute the backward KL in float64.
@@
(a) Resetting the optimizer state only.
@@
(a) Perplexity of completions when distilling with the KLD objective.
(b) Perplexity of completions when distilling with the MSE objective.
(c) Perplexity of completions when distilling with the TVD objective.
Figure 18: Conditional perplexity. Perplexity of the completions using GPT-2 large, excluding the prompt. SDTT with TVD performs worse. The final student distilled with KLD matches GPT-2 with nucleus sampling. Ground-truth continuations have a perplexity $\approx 13.11$.
(a) Distillation with the MSE loss.
(b) Distillation with the TVD loss.
Figure 19: Diversity of conditional generation (small scale). We measure the trade-off between quality and diversity using Self-BLEU (Zhu et al., 2018). Deterministic sampling yields a score of 1. The diversity minimally decreases after distillation.
(a) Small (169M.
(b) Medium (424M).
(c) Large (863M).
(d) 1.3B.
(e) 3B.
(f) 8B.
Figure 20: Additional latency experiments with a batch size of 8.
(a) Small (169M.
(b) Medium (424M).
(c) Large (863M).
(d) 1.3B.
(e) 3B.
(f) 8B.
Figure 21: Additional latency experiments with a batch size of 4.
Figure 22: Perplexity vs wall-time latency (in seconds) for small models. We use 16, 32, 64, 128 ans 256 decoding step for the diffusion models.
@@ Appendix B Additional details on the divergence measures
Model size small medium large 1.3B 3B 8B
# params 169M 424M 863M 1.3B 3B 8B
Num Layers 12 24 24 24 26 40
Embedding dim. 768 1024 1536 2048 3072 4096
Num. heads 12 16 16 32 32 32
Table 2: Hyperparameters of the diffusion models at different scales. All models use RoPE positional encoding (Su et al., 2023).
@@ B.3 Mean-Squared Error
B.4 $\chi^{2}$ divergence
The $\chi^{2}$ divergence can be used to compare two probability distributions. For two discrete probability distributions $p$ and $q$ defined on the same sample space $\Omega$m the $\chi^{2}$ divergence is computed as:
${d_{\chi^{2}}(p,q)=\sum_{x\in\Omega}q(x)\left(\frac{p(x)}{q(x)}-1\right)^{2}=\sum_{x\in\Omega}\frac{1}{q(x)}\left(p(x)-q(x)\right)^{2}.}$ (12)
As such, we see that the $\chi^{2}$ divergence is related to the MSE. Note that when using the MSE for distillation, we penalize the error in log space, while the $\chi^{2}$ penalizes error in probability space. Additionally, the MSE uses a uniform weight factor $\frac{1}{|\Omega|}$ for each term of the sum, while the $\chi^{2}$ divergence uses a weight of $\frac{1}{q(x)}$.
@@ Appendix D Text examples
We include non-cherry picked text generated from the small distilled model with KLD loss from the last round of distillation via unconditional sampling with varying number of steps. We show the first 512 tokens to so that the text fits on one page. Remember that those models are small and not fine-tuned for text quality. They can also start generating in the middle of sentences, since they are trained on a concatenated corpus of documents.